# Motivic classes of stacks in finite characteristic and applications to stacks of Higgs bundles

### v1

@@ 2.4.1. Bilinear form
Let $\mathcal{Z}$ be a $k$-stack of finite type. If $\mathcal{X}$ and $\mathcal{Y}$ are of finite type over $\mathcal{Z}$, we set
$([\mathcal{X}\to\mathcal{Z}]|[\mathcal{Y}\to\mathcal{Z}])\coloneqq[\mathcal{X}\times_{\mathcal{Z}}\mathcal{Y}].$
We get a symmetric bilinear form $\operatorname{Mot}(\mathcal{Z})\otimes\operatorname{Mot}(\mathcal{Z})\to\operatorname{Mot}(k)$ by extending this by bilinearity. This can also be extended to a symmetric form $\overline{\operatorname{Mot}}(\mathcal{Z})\otimes\overline{\operatorname{Mot}}(\mathcal{Z})\to\overline{\operatorname{Mot}}(k)$ by continuity.
Let $\mathcal{Z}$ be a $k$-stack locally of finite type. Given $A\in\overline{\operatorname{Mot}}^{fin}(\mathcal{Z})$, $B\in\overline{\operatorname{Mot}}(\mathcal{Z})$, we write $A=j_{!}A_{\mathcal{V}}$, where $\mathcal{V}\in\operatorname{Opf}(\mathcal{Z})$ with open immersion $j\colon\mathcal{V}\to\mathcal{X}$. Let $B$ be given by an inverse system $\mathcal{U}\mapsto B_{\mathcal{U}}$, where $\mathcal{U}$ ranges over $\operatorname{Opf}(\mathcal{Z})$. Set $(A|B)\coloneqq(A_{\mathcal{V}}|B_{\mathcal{V}})$, which is dependent of the choice of $\mathcal{V}$. Now we have a continuous bilinear form
$(\bullet|\bullet)\colon\overline{\operatorname{Mot}}^{fin}(\mathcal{Z})\otimes\overline{\operatorname{Mot}}(\mathcal{Z})\to\overline{\operatorname{Mot}}(k).$
We will get a symmetric bilinear form if we restrict to $\overline{\operatorname{Mot}}^{fin}(\mathcal{Z})\otimes\overline{\operatorname{Mot}}^{fin}(\mathcal{Z})$.
Remark 2.21.
Let us call the group $\operatorname{Mot}(\mathcal{X})$ defined in [FSS18, Section 2.1] without surjective and universally injective relation $\operatorname{Mot}^{naive}(\mathcal{X})$. We can check whether the motivic functions are equal in $\operatorname{Mot}^{naive}(\mathcal{X})$ fiberwise, see [FSS18, Proposition 2.6.1].
Currently, we don’t know whether the equality of motivic functions can be checked fiberwise either in $\operatorname{Mot}(\mathcal{X})$ or in the completion $\overline{\operatorname{Mot}}(\mathcal{X})$, but we can first check an identity in $\operatorname{Mot}^{naive}(\mathcal{X})$ and then apply the homomorphism $\operatorname{Mot}^{naive}(\mathcal{X})\to\operatorname{Mot}(\mathcal{X})$. See the last paragraph of Section 5.3 below for an example.
@@ 2.5. Motivic classes over the perfection of the field
Finally we want to give a theorem comparing the motivic classes over a field $k$ and its perfection.
@@
We have an isomorphism of rings
$\operatorname{Mot}(k)\xrightarrow{\sim}\operatorname{Mot}(k^{perf}).$
@@
Now it suffices to prove
$\operatorname{Mot}(k)\simeq\operatorname{Mot}(k^{1/p}).$
By Theorem 2.6, it suffices to prove that
$\operatorname{Mot}_{sch}(k)\simeq\operatorname{Mot}_{sch}(k^{1/p}).$
There is a natural homomorphism
$\operatorname{Mot}_{sch}(k)\to\operatorname{Mot}_{sch}(k^{1/p})$
defined by the rule $Y]\mapsto[Y\times_{k}k^{1/p}]$ for any scheme $Y$ over $k$.
The inverse homomorphism is the pushforward along the finite morphism
$\operatorname{Spec}k^{1/p}\to\operatorname{Spec}k$
from the previous section.
Once the homomorphisms are defined, we need to check that the compositions are identity.
One composition takes a $k^{1/p}$-scheme $X$ to its Frobenius twist $X^{(p)}\coloneqq X\times_{k}k^{1/p}$ over $k$. We have a morphism $X^{(p)}\to X$ that is surjective and universally injective because it is a base change of $\operatorname{Spec}k^{1/p}\to\operatorname{Spec}k$.
The other composition takes a $k$-scheme $Y$ to $Y\times_{k}k^{1/p}$ viewed as a $k$-scheme. This composition is the identity for a similar reason. ∎
@@ 2.6. The universal $\lambda$-ring quotient
In this section, we recall the definition of the universal quotient of $\overline{\operatorname{Mot}}(k)$ that is a $\lambda$-ring, see [LL04] and [FSS24, Section 8] for more details.
Let $R$ be a pre-$\lambda$-ring (see Definition 2.15), and write $\lambda\colon R\to(1+zR[[z]])^{\times}$ as $\lambda=\sum_{i=0}^{\infty}\lambda_{i}z^{i}$.
Recall that a $\lambda$-ring is a pre-$\lambda$-ring $R$ such that $\lambda(1)=1+z$ and for all $x,y\in R$ and $m,n\in\mathbb{Z}_{\geq 0}$, $\lambda_{n}(xy)$ and $\lambda_{n}(\lambda_{m}(x))$ can be expressed in terms of $\lambda_{i}(x)$ and $\lambda_{j}(y)$ using the ring operations in a standard way (see [SGA6, Exposé V, Définition 2.4, (2.4.1)] and [Knu73, Chapter I., Section 1]).
@@
We don’t know whether the following composition is injective:
@@ 3. Counting measure for the ring of motivic classes
3.1. Counting measure for the ring of motivic classes
In this section we want to define a counting measure for the stacks over finite fields.
First we give the definition of the volume of a groupoid.
@@ 3.1. Counting measure for the ring of motivic classes
Now we denote by $|\mathcal{X}(\mathbb{F}_{q})|$ the volume of the groupoid $\mathcal{X}(\mathbb{F}_{q})$. Note that for a scheme $X$, $|X(\mathbb{F}_{q})|$ coincides with the number of $\mathbb{F}_{q}$-points of $X$.
Proposition 3.2.
@@
Proof.
@@
By Theorem 2.6, we have
$\operatorname{Mot}_{sch}(\mathbb{F}_{q})[\mathbb{L}^{-1},(\mathbb{L}^{i}-1)^{-1}|i>0]\cong\operatorname{Mot}(\mathbb{F}_{q}).$
By [Kre99, Lemma 3.5.1, Propositions 3.5.6 and 3.5.9], it suffices to show that for a scheme $X$,
$|X/\operatorname{GL}(n)|=\frac{\#(X)}{\#(\operatorname{GL}(n))}.$
We are done by [Beh93, Lemma 2.5.1]. The uniqueness is obvious. ∎
This ring homomorphism is called a counting measure.
Now we denote by $\overline{\operatorname{Mot}}_{fin-type}(\mathbb{F}_{q})$ the image of $\operatorname{Mot}(\mathbb{F}_{q})$ in $\overline{\operatorname{Mot}}(\mathbb{F}_{q})$. We want to pass from pre-$\lambda$-rings to $\lambda$-rings (see Section 2.6). We denote by $\overline{\operatorname{Mot}}_{fin-type}^{\lambda-ring}(\mathbb{F}_{q})$ the image of the composition
Theorem 3.3.
We have a unique ring homomorphism
$\colon\overline{\operatorname{Mot}}_{fin-type}^{\lambda-ring}(\mathbb{F}_{q})\to\mathbb{Q}$
Since $\overline{\operatorname{Mot}}_{fin-type}^{\lambda-ring}(\mathbb{F}_{q})$ is a quotient of $\operatorname{Mot}(\mathbb{F}_{q})$, in view of Proposition 3.2, the theorem is equivalent to the statement that for any $\mathcal{X}$ of finite type, $|\mathcal{X}(\mathbb{F}_{q})|$ depends only on the image of $\mathcal{X}$ in $\overline{\operatorname{Mot}}^{\lambda-ring}(\mathbb{F}_{q})$.
@@
Our proof is inspired by [BD07, Section 5], where for a subring $\overline{\operatorname{Mot}}_{conv}(\mathbb{F}_{q})$ of $\overline{\operatorname{Mot}}(\mathbb{F}_{q})$ consisting of convergent elements, the authors defined a counting measure
@@
To prove the above theorem, first we need to define the following $\lambda$-ring $R_{q}$.
@@ 3.2. $\lambda$-ring $R_{q}$
We first give the definition of Weil number: $\alpha\in\overline{\mathbb{Q}}$ is called a $q$-Weil number of weight $i$, if for all embeddings $\sigma\colon\overline{\mathbb{Q}}\hookrightarrow\mathbb{C}$, we have $|\sigma(\alpha)|=q^{i/2}$.
We define the following abelian category
@@
The isomorphism classes of the simple objects in $J_{q}$ are $E_{\lambda}=(\overline{\mathbb{Q}}_{\ell},\lambda)$, where $\lambda\in\overline{\mathbb{Q}}_{\ell}=\operatorname{End}(\overline{\mathbb{Q}}_{\ell})$ is such that $\lambda$ is a $q$-Weil number. Thus, $R_{q}$ as a $\mathbb{Q}$-vector space has a basis $E_{\lambda}]$ and the product is $E_{\lambda}][E_{\mu}]=[E_{\lambda\mu}]$.
Since $\mathbb{Q}\subset R_{q}$, by [FSS24, Section 8.1], we can define the $\lambda$-ring structure by specifying the $\psi$-operations. We define the $\psi$-operations to be the homomorphisms $\psi_{n}\colon R_{q}\to R_{q}$ such that for the basis $E_{\lambda}]\in R_{q}$, $\psi_{n}([E_{\lambda}])=[E_{\lambda^{n}}]$. Now we have the following properties $\psi_{0}=1,\psi_{1}=\operatorname{Id}_{R_{q}}$, and for all $n,m\in\mathbb{Z}\geq 0$ we have $\psi_{nm}=\psi_{n}\circ\psi_{m}$. This gives us a well-defined $\lambda$-ring structure.
@@
For $V_{\pm}\in J_{q}$ we have
$\lambda([V_{+}]-[V_{-}])=\left(\sum_{i=0}^{\dim V_{+}}[\wedge^{i}V_{+}]z^{i}\right)\left(\sum_{i=0}^{\infty}[Sym^{i}V_{-}](-z)^{i}\right).$
@@
$\lambda([V_{+}])=\sum_{i=0}^{\dim V_{+}}[\wedge^{i}V_{+}]z^{i}.$
For an exact sequence $0\to V_{1}\to V_{+}\to V_{2}\to 0$, there is a filtration
$0=F^{0}\subseteq F^{1}\subseteq F^{2}\subseteq\cdots\subseteq F^{i}=\wedge^{i}V_{+},$
such that $F^{j}/F^{j-1}\simeq\wedge^{j}V_{1}\otimes\wedge^{i-j}V_{2}$. Thus,
$\wedge^{i}V_{+}]=\bigoplus_{j=0}^{i}[\wedge^{j}V_{1}]\cdot[\wedge^{i-j}V_{2}].$
It follows that it is enough to check the identity when $V_{+}=E_{\lambda}$. By [FSS24, Lemma 8.1.1],
$\lambda([E_{\lambda}])=1+z[E_{\lambda}]=\sum_{i=0}^{1}[\wedge^{i}E_{\lambda}]z^{i}.$
@@
$\lambda(-[E_{\lambda}])=\frac{1}{1+z[E_{\lambda}]}=\sum_{i=0}^{\infty}[Sym^{i}E_{\lambda}](-z)^{i},$
@@
$\lambda(-[V_{-}])=\sum_{i=0}^{\infty}[Sym^{i}V_{-}](-z)^{i}.$
We can re-write the above formula using super objects. Let $V=V_{-}\oplus V_{+}$ be a super object of $J_{q}$, we define $V]\coloneqq[V_{+}]-[V_{-}]$. We have
$Sym^{i}V\coloneqq\bigoplus_{j=0}^{i}\left(Sym^{j}V_{+}\otimes\wedge^{i-j}V_{-}\right).$
@@
For $V=V_{-}\oplus V_{+}$ a super object of $J_{q}$, we have
$\lambda(-[V])=\sum_{i=0}^{\infty}[Sym^{i}V](-z)^{i}.$
@@
By Lemma 3.5,
$\lambda(-[V])$ $=\lambda([V_{-}]-[V_{+}])$
$=\left(\sum_{i=0}^{\infty}[Sym^{i}V_{+}](-z)^{i}\right)\left(\sum_{i=0}^{\dim V_{+}}[\wedge^{i}V_{-}]z^{i}\right)$
$=\sum_{i=0}^{\infty}[Sym^{i}V](-z)^{i}.$
The last equality holds because $Sym^{i}V]=\sum_{j}(-1)^{i-j}[Sym^{j}V_{+}][\wedge^{i-j}V_{-}]$. ∎
@@
Now we construct the étale realization of motivic classes and give the proof of Theorem 3.3.
We define $\overline{X}\coloneqq X\otimes_{\operatorname{Spec}\mathbb{F}_{q}}\operatorname{Spec}\overline{\mathbb{F}}_{q}$ where $\overline{\mathbb{F}}_{q}$ is a fixed algebraic closure of $\mathbb{F}_{q}$ and denote by $F_{q}$ the Frobenius action on the $\ell$-adic cohomology.
@@
sending the class $[X]$ of a quasi-projective scheme $X$ to $\sum_{i}(-1)^{i}[H^{i}_{c}(\overline{X},\overline{\mathbb{Q}}_{\ell}),F_{q}]\in R_{q}.$
@@
Recall the definition of $\operatorname{Mot}_{sch}(k)$ in Definition 2.3. We claim that there is a ring homomorphism $H^{\ast}\colon\operatorname{Mot}_{sch}(\mathbb{F}_{q})\to R_{q}$ such that
@@
    $H^{\ast}([X])=H^{\ast}([Y])+H^{\ast}([X-Y])$ where $Y$ is a closed subscheme of $X$ over $\mathbb{F}_{q}$;
@@
    $H^{\ast}([X])=H^{\ast}([Y])$ if there is a surjective and universally injective morphism $X\to Y$ of schemes over $\mathbb{F}_{q}$;
@@
    $H^{\ast}([X\times Y])=H^{\ast}([X])\cdot H^{\ast}([Y])$ where $X$, $Y$ are schemes over $\mathbb{F}_{q}$.
The first relation holds by the long exact sequence of the cohomology, see [SGA4, Exposé XVII, (5.1.16.3)]. For the second relation, by Lemma 3.8, we know there is a stratification of universal homeomorphisms. Then by [SP18, Tag03SI], we are done. Finally the third relation holds by Künneth theorem, see [SGA4, Exposé XVII, Théorème 5.4.3] and [Kat94, Page 24].
Note that $H^{\ast}(\mathbb{L})=[\overline{\mathbb{Q}}_{\ell},q]$. Since $L\coloneqq[\overline{\mathbb{Q}}_{\ell},q]$ is invertible in $R_{q}$, we can extend $H^{\ast}$ to $\operatorname{Mot}_{sch}(\mathbb{F}_{q})[\mathbb{L}^{-1}]$. We know that the absolute values of the eigenvalues of $F_{q}$ acting on $H^{i}_{c}(\overline{X},\overline{\mathbb{Q}}_{\ell})$ are bounded above by $q^{i/2}$ by [Del80, Corollaire 3.3.4]. Also $H^{i}_{c}(\overline{X},\overline{\mathbb{Q}}_{\ell})]=0$ for $i>2\dim X$, then we have $H^{\ast}([X])\in\bigoplus_{i\leq 2\dim X}R_{q,i}$.

### v2

@@ 2.4.1. Bilinear form
Let $\mathcal{Y}$ be a $k$-stack of finite type. If $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ are of finite type over $\mathcal{Y}$, we set
$([\mathcal{X}_{1}\to\mathcal{Y}]|[\mathcal{X}_{2}\to\mathcal{Y}])\coloneqq[\mathcal{X}_{1}\times_{\mathcal{Y}}\mathcal{X}_{2}].$
We get a symmetric bilinear form $\operatorname{Mot}(\mathcal{Y})\otimes\operatorname{Mot}(\mathcal{Y})\to\operatorname{Mot}(k)$ by extending this by bilinearity. This can also be extended to a symmetric form $\overline{\operatorname{Mot}}(\mathcal{Y})\otimes\overline{\operatorname{Mot}}(\mathcal{Y})\to\overline{\operatorname{Mot}}(k)$ by continuity.
Let $\mathcal{Y}$ be a $k$-stack locally of finite type. Given $A\in\overline{\operatorname{Mot}}^{fin}(\mathcal{Y})$, $B\in\overline{\operatorname{Mot}}(\mathcal{Y})$, we write $A=j_{!}A_{\mathcal{V}}$, where $\mathcal{V}\in\operatorname{Opf}(\mathcal{Y})$ with open immersion $j\colon\mathcal{V}\to\mathcal{Y}$. Let $B$ be given by an inverse system $\mathcal{U}\mapsto B_{\mathcal{U}}$, where $\mathcal{U}$ ranges over $\operatorname{Opf}(\mathcal{Y})$. Set $(A|B)\coloneqq(A_{\mathcal{V}}|B_{\mathcal{V}})$, which is independent of the choice of $\mathcal{V}$. Now we have a continuous bilinear form
$(\bullet|\bullet)\colon\overline{\operatorname{Mot}}^{fin}(\mathcal{Y})\otimes\overline{\operatorname{Mot}}(\mathcal{Y})\to\overline{\operatorname{Mot}}(k).$
We will get a symmetric bilinear form if we restrict to $\overline{\operatorname{Mot}}^{fin}(\mathcal{Y})\otimes\overline{\operatorname{Mot}}^{fin}(\mathcal{Y})$.
@@ 2.5. Motivic classes over the perfection of the field
We want to give a theorem comparing the motivic classes over a field $k$ and its perfection.
@@
Let $f:\operatorname{Spec}k^{perf}\to\operatorname{Spec}k$ be the natural morphism. Then the morphism
$f^{\ast}:\operatorname{Mot}(k)\xrightarrow{\sim}\operatorname{Mot}(k^{perf})$
is an isomorphism of rings.
@@
We have a natural morphism $g:\operatorname{Spec}k^{1/p}\to\operatorname{Spec}k$. Now it suffices to prove that the induced morphisms $g^{\ast}$ and $g_{\ast}$ are inverse to each other
$g^{\ast}:\operatorname{Mot}(k)\xrightarrow{\sim}\operatorname{Mot}(k^{1/p}):g_{\ast}.$
By Theorem 2.6, it suffices to prove a similar statement with $\operatorname{Mot}(k)$ replaced by $\operatorname{Mot}_{sch}(k)$. One composition takes the class of a $k^{1/p}$-scheme $X$ to the class of its Frobenius twist $X^{(p)}\coloneqq X\times_{k}k^{1/p}$ over $k$. We have a morphism $X^{(p)}\to X$ that is surjective and universally injective because it is a base change of $\operatorname{Spec}k^{1/p}\to\operatorname{Spec}k$.
The other composition takes the class of a $k$-scheme $Y$ to the class of $Y\times_{k}k^{1/p}$ viewed as a $k$-scheme. This composition is the identity for a similar reason. ∎
@@ 2.6. The universal $\lambda$-ring quotient
We recall the definition of the universal quotient of $\overline{\operatorname{Mot}}(k)$ that is a $\lambda$-ring, see [LL04] and [FSS24, Section 8] for more details.
Let $R$ be a pre-$\lambda$-ring (see Definition 2.15), and write $\lambda\colon R\to(1+zR[[z]])^{\times}$ as $\lambda=\sum_{i=0}^{\infty}\lambda_{i}z^{i}$. Recall that $R$ is a $\lambda$-ring if $\lambda(1)=1+z$ and for all $x,y\in R$ and $m,n\in\mathbb{Z}_{\geq 0}$, $\lambda_{n}(xy)$ and $\lambda_{n}(\lambda_{m}(x))$ can be expressed in terms of $\lambda_{i}(x)$ and $\lambda_{j}(y)$ using the ring operations in a standard way (see [SGA6, Exposé V, Définition 2.4, (2.4.1)] and [Knu73, Chapter I., Section 1]).
@@
We don’t know whether the following two morphisms are injective:
@@ 3. Counting measure for the ring of motivic classes
In this section we define a counting measure for the stacks of finite type over finite fields and show that the volume of the groupoid only depends on the image in the universal $\lambda$-ring quotient.
3.1. Counting measure for the universal $\lambda$-ring quotient
Recall the definition of the volume of a groupoid.
@@ 3.1. Counting measure for the universal $\lambda$-ring quotient
Fix a prime power $q$. If $\mathcal{X}$ is a stack of finite type over the finite field $\mathbb{F}_{q}$, we denote by $|\mathcal{X}(\mathbb{F}_{q})|$ the volume of the groupoid $\mathcal{X}(\mathbb{F}_{q})$. Note that for a scheme $X$, $|X(\mathbb{F}_{q})|$ coincides with the number of $\mathbb{F}_{q}$-points of $X$.
Theorem 3.2.
  1. (i)
@@
  2. (ii)
    For any $A\in\operatorname{Mot}(\mathbb{F}_{q})$, the counting measure #(A)\#(A) depends only on the image of $A$ in $\overline{\operatorname{Mot}}^{\lambda-ring}(\mathbb{F}_{q})$.
Proof of part (i).
@@
Composing Theorem 2.6 and [Kre99, Lemma 3.5.1, Propositions 3.5.6 and 3.5.9], it suffices to show that for a scheme $X$,
$|\left(X/\operatorname{GL}(n)\right)(\mathbb{F}_{q})|=\frac{\#([X])}{\#([\operatorname{GL}(n)])}.$
By the proof of Theorem 2.6 and the definition of #\#,
$([X/\operatorname{GL}(n)])=\frac{\#([X])}{\#([\operatorname{GL}(n)])}.$
Applying [Beh93, Lemma 2.5.1], we finally have
$|\left(X/\operatorname{GL}(n)\right)(\mathbb{F}_{q})|=\frac{\#([X])}{\#([\operatorname{GL}(n)])}=\#([X/\operatorname{GL}(n)]).$
The uniqueness is obvious. ∎
This ring homomorphism is called the counting measure.
@@
Our proof is inspired by [BD07, Section 5], where for a subring $\overline{\operatorname{Mot}}_{conv}(\mathbb{F}_{q})$ of $\overline{\operatorname{Mot}}(\mathbb{F}_{q})$ consisting of the so-called convergent elements, the authors defined a counting measure
@@
To prove the second part of the above theorem, first we need to define the following $\lambda$-ring $R_{q}$.
@@ 3.2. $\lambda$-ring $R_{q}$
We first recall the definition of Weil number: $\alpha\in\overline{\mathbb{Q}}$ is called a $q$-Weil number of weight $i$, if for all embeddings $\sigma\colon\overline{\mathbb{Q}}\hookrightarrow\mathbb{C}$, we have $|\sigma(\alpha)|=q^{i/2}$.
We define the following abelian symmetric monoidal category
@@
The isomorphism classes of the simple objects in $J_{q}$ are $\mathcal{E}_{\lambda}\coloneqq(\overline{\mathbb{Q}}_{\ell},\lambda)$, where $\lambda\in\overline{\mathbb{Q}}_{\ell}=\operatorname{End}(\overline{\mathbb{Q}}_{\ell})$ is such that $\lambda$ is a $q$-Weil number. Thus, $R_{q}$ as a $\mathbb{Q}$-vector space has a basis $\mathcal{E}_{\lambda}]$ and the product is $\mathcal{E}_{\lambda}][\mathcal{E}_{\mu}]=[\mathcal{E}_{\lambda\mu}]$.
Since $\mathbb{Q}\subset R_{q}$, by [FSS24, Section 8.1], we can define the $\lambda$-ring structure by specifying the $\psi$-operations. We define the $\psi$-operations to be the homomorphisms $\psi_{n}\colon R_{q}\to R_{q}$ such that for the basis $\mathcal{E}_{\lambda}]\in R_{q}$, $\psi_{n}([\mathcal{E}_{\lambda}])=[\mathcal{E}_{\lambda^{n}}]$. Now we have the following properties $\psi_{0}=1,\psi_{1}=\operatorname{Id}_{R_{q}}$, and for all $n,m\in\mathbb{Z}_{\geq 0}$ we have $\psi_{nm}=\psi_{n}\circ\psi_{m}$. This gives us a well-defined $\lambda$-ring structure.
@@
For $\mathcal{V}_{\pm}\in J_{q}$ we have
$\lambda([\mathcal{V}_{+}]-[\mathcal{V}_{-}])=\left(\sum_{i=0}^{\dim\mathcal{V}_{+}}[\wedge^{i}\mathcal{V}_{+}]z^{i}\right)\left(\sum_{i=0}^{\infty}[Sym^{i}\mathcal{V}_{-}](-z)^{i}\right).$
@@
$\lambda([\mathcal{V}_{+}])=\sum_{i=0}^{\dim\mathcal{V}_{+}}[\wedge^{i}\mathcal{V}_{+}]z^{i}.$
For an exact sequence $0\to\mathcal{V}_{1}\to\mathcal{V}_{+}\to\mathcal{V}_{2}\to 0$, there is a filtration
$0=\mathcal{F}^{0}\subseteq\mathcal{F}^{1}\subseteq\mathcal{F}^{2}\subseteq\cdots\subseteq\mathcal{F}^{i}=\wedge^{i}\mathcal{V}_{+},$
such that $\mathcal{F}^{j}/\mathcal{F}^{j-1}\simeq\wedge^{j}\mathcal{V}_{1}\otimes\wedge^{i-j}\mathcal{V}_{2}$. Thus,
$\wedge^{i}\mathcal{V}_{+}]=\bigoplus_{j=0}^{i}[\wedge^{j}\mathcal{V}_{1}]\cdot[\wedge^{i-j}\mathcal{V}_{2}].$
It follows that it is enough to check the identity when $\mathcal{V}_{+}=\mathcal{E}_{\lambda}$. By [FSS24, Lemma 8.1.1],
$\lambda([\mathcal{E}_{\lambda}])=1+z[\mathcal{E}_{\lambda}]=\sum_{i=0}^{1}[\wedge^{i}\mathcal{E}_{\lambda}]z^{i}.$
@@
$\lambda(-[\mathcal{E}_{\lambda}])=\frac{1}{1+z[\mathcal{E}_{\lambda}]}=\sum_{i=0}^{\infty}[Sym^{i}\mathcal{E}_{\lambda}](-z)^{i},$
@@
$\lambda(-[\mathcal{V}_{-}])=\sum_{i=0}^{\infty}[Sym^{i}\mathcal{V}_{-}](-z)^{i}.$
We can re-write the above formula using super objects. Let $\mathcal{V}=\mathcal{V}_{-}\oplus\mathcal{V}_{+}$ be a super object of $J_{q}$, we define $\mathcal{V}]\coloneqq[\mathcal{V}_{+}]-[\mathcal{V}_{-}]\in R_{q}$. We have
$Sym^{i}\mathcal{V}\coloneqq\bigoplus_{j=0}^{i}\left(Sym^{j}\mathcal{V}_{+}\otimes\wedge^{i-j}\mathcal{V}_{-}\right).$
@@
Let $\mathcal{V}$ be a super object of $J_{q}$, we have
$\lambda(-[\mathcal{V}])=\sum_{i=0}^{\infty}[Sym^{i}\mathcal{V}](-z)^{i}.$
@@
Write $\mathcal{V}=\mathcal{V}_{-}\oplus\mathcal{V}_{+}$. By Lemma 3.4,
$\lambda(-[\mathcal{V}])$ $=\lambda([\mathcal{V}_{-}]-[\mathcal{V}_{+}])$
$=\left(\sum_{i=0}^{\infty}[Sym^{i}\mathcal{V}_{+}](-z)^{i}\right)\left(\sum_{i=0}^{\dim\mathcal{V}_{+}}[\wedge^{i}\mathcal{V}_{-}]z^{i}\right)$
$=\sum_{i=0}^{\infty}[Sym^{i}\mathcal{V}](-z)^{i}.$
The last equality holds because $Sym^{i}\mathcal{V}]=\sum_{j}(-1)^{i-j}[Sym^{j}\mathcal{V}_{+}][\wedge^{i-j}\mathcal{V}_{-}]$. ∎
@@
Now we construct the étale realization of motivic classes and give the proof of the second part of Theorem 3.2.
Let $X$ be a reduced quasi-projective scheme over $\mathbb{F}_{q}$. We define
$\overline{X}\coloneqq X\times_{\mathbb{F}_{q}}\operatorname{Spec}\overline{\mathbb{F}}_{q}$
where $\overline{\mathbb{F}}_{q}$ is a fixed algebraic closure of $\mathbb{F}_{q}$ and denote by $F_{q}$ the Frobenius action on the $\ell$-adic cohomology.
@@
sending the class $[X]$ of a reduced quasi-projective scheme $X$ to the following class:
$\sum_{i}(-1)^{i}[H^{i}_{c}(\overline{X},\overline{\mathbb{Q}}_{\ell}),F_{q}]\in R_{q}.$
@@
Step 1. Recall the definition of $\operatorname{Mot}_{sch}(k)$ in Definition 2.3. We claim that there is a ring homomorphism $H^{\ast}\colon\operatorname{Mot}_{sch}(\mathbb{F}_{q})\to R_{q}$ such that
@@
    $H^{\ast}([X])=H^{\ast}([Y])+H^{\ast}([X-Y])$ where $Y$ is a closed subscheme of $X$ over $\mathbb{F}_{q}$.
@@
    $H^{\ast}([X])=H^{\ast}([Y])$ if there is a surjective and universally injective morphism $X\to Y$ of schemes over $\mathbb{F}_{q}$.
@@
    $H^{\ast}([X\times_{\mathbb{F}_{q}}Y])=H^{\ast}([X])\cdot H^{\ast}([Y])$ where $X$, $Y$ are schemes over $\mathbb{F}_{q}$.
The first relation holds by the long exact sequence of the cohomology, see [SGA4, Exposé XVII, (5.1.16.3)]. For the second relation, by Lemma 3.7, we know there is a stratification by universal homeomorphisms. Then by [SP18, Tag03SI], a universal homeomorphism induces an isomorphism of étale sites, we are done. Finally the third relation holds by Künneth theorem, see [SGA4, Exposé XVII, Théorème 5.4.3] and [Kat94, Page 24].
Step 2. Note that $H^{\ast}(\mathbb{L})=[\overline{\mathbb{Q}}_{\ell},q]$. Since $\mathcal{L}\coloneqq[\overline{\mathbb{Q}}_{\ell},q]$ is invertible in $R_{q}$, we can extend $H^{\ast}$ to $\operatorname{Mot}_{sch}(\mathbb{F}_{q})[\mathbb{L}^{-1}]$. We know that the eigenvalues of the Frobenius acting on $i$-th cohomology group are $q$-Weil numbers of weights bounded by $i$ by [Del80, Corollaire 3.3.4]. Also $H^{i}_{c}(\overline{X},\overline{\mathbb{Q}}_{\ell})=0$ for $i>2\dim X$, then we have $H^{\ast}([X])\in\bigoplus_{i\leq 2\dim X}R_{q,i}$.